if a rectangle has width of 5 and diagonal with lenght of 13, what is the area of the rectangle?
Answer by nerdybill(7384)

if a rectangle has width of 5 and diagonal with lenght of 13, what is the area of the rectangle?
.
find the length by applying Pythagorean theorem:
Let x = length
then
x^2 + 5^2 = 13^2
x^2 = 13^2 - 5^2
x^2 = 169 - 25
x^2 = 144
x = 12
.
area = 5*12 = 60 square units